OpenStudy (anonymous):

90% of people can't solve this??? Maximizing income?? A manufacturer is producing two types of unites. Each unit Q costs 9$ for parts and $15 for labor and each uniot R costs $6 for parts and $20 for labor. The manufacturer's budget is $810 for parts and $1800 for labor. If the income per unit is $150 for Q and $175 for R, how many units of each should be manufactured to maximise income?

OpenStudy (amistre64):

setup your constraints ....

OpenStudy (anonymous):

How? @amistre64

OpenStudy (amistre64):

this makes the ost sense to me Q = 9p + 15l R = 6p + 20l parts: 9Q + 6R = 810 labor: 15Q + 20R = 1800 income: 150Q + 175R

OpenStudy (amistre64):

not saying its right, but shouldnt these conditions hold?

OpenStudy (anonymous):

Slowly starting to piece something together here :/

OpenStudy (anonymous):

That's starting to make a lot more sense

OpenStudy (perl):

I googled this question and got "A manufacturer is producing 100 total units. Each unit Q costs $9 for parts and $15 for labor, and each unit R costs $6 and $20 for labor. The manufacturer’s budget is $810 for parts and $1800 for labor. If the income per unit $150 for Q and $175 for R, how many units of each should be manufactured to maximize income?"

OpenStudy (amistre64):

0Q and 0R of course is the minimum income ... nothing made, nothing sold lets just sell Q 9Q = 810 ; Q=90 <--- this limits out our budget sooo 15Q = 1500 ; Q=100 150(90) is our income for just selling Q within our budget 13500 ------------------------------ maybe just R? 6R = 810 ; R=135 20R = 1500 ; R=75 <--- this limits out our budget sooo 175(75) = 13125 -------------------------------- spose we make 1R how many Qs can we make within budget? 9Q + 6 = 810 ; Q = 89 <---- limited to 89 to be in budget 15Q + 20= 1500 ; Q = 98 ------------------------------- income is: 150(89) + 175(1) = 13525 you see how this is working?

OpenStudy (amistre64):

1800, not 1500 .... had some typos in mine

OpenStudy (amistre64):

why Q+R = 100?

OpenStudy (amistre64):

im getting 104 ... when i find the max values

OpenStudy (perl):

I may have found a variation on the problem.

OpenStudy (amistre64):

Q = 9p + 15l R = 6p + 20l parts: 9(9p + 15l) + 6(6p + 20l) = 810 labor: 15(9p + 15l) + 20(6p + 20l) = 1800 income: 150Q + 175R p = 35/6 l = 1/2 Q = 9(35/6) + 15(1/2) = 60 R = 6(35/6) + 20(1/2) = 45 150(61) + 175(43) = 16875

OpenStudy (amistre64):

ok, my notepad version of that has some sidereal calculation lol 60 and 45 is what i get as a max, and i tried to alter it to test

OpenStudy (amistre64):

150(60) + 175(45) = 16875 150(61) + 175(43) = 16675 a decrease

OpenStudy (anonymous):

Thanks so much

OpenStudy (perl):

did you do something like this http://prntscr.com/72qqwn

OpenStudy (perl):

plug in the vertices

OpenStudy (amistre64):

in a round about way yes

OpenStudy (perl):

plugging in the corner points (0,0), (0,90), (60,45), and (90,0) into the objective function income: 150x + 175y , it is maximized by (60,45)

OpenStudy (perl):

So you can solve it two ways, looks like amistre did it by direct substitution, which makes sense since if you maximize the budget for each, you will get the most stuff to sell and therefore the most income. As a linear programming problem: Objective function: 150Q + 175R Constraints: 9Q + 6R <= 810 15Q + 20R <=1800 Q>=0 R>=0 The last two constraints force us to look only in the first quadrant and you will graph the lines 9Q + 6R = 810 15Q + 20R =1800 , find the vertices, plug them into the objective function